# Query Optimization for Parametric Knowledge Refinement in Retrieval-Augmented Large Language Models

### v3

@@ Abstract
Youan Congthanks: Equal contribution.   Pritom Saha Akash11footnotemark: 1   Cheng Wang   Kevin Chen-Chuan Chang University of Illinois Urbana-Champaign, USA {youanc2, pakash2, chengw4, kcchang}@illinois.edu
@@ 1 Introduction
The field of natural language processing (NLP) has witnessed transformative advancements in recent years, largely driven by the advent of Large Language Models (LLMs). These models, trained on vast corpora, have demonstrated exceptional capabilities in understanding human text and generating high-quality responses Kaplan et al. (2020); Clark et al. (2022). They have also proven practical and scalable for various downstream NLP tasks, such as conversational response generation, text summarization, and content recommendation, even in few-shot or zero-shot settings Wu et al. (2023). Despite their strengths, a key limitation of LLMs lies in their reliance on static training data, which makes them struggle with dynamic or less commonly known information outside their initial training scope. This limitation often leads to outdated, inaccurate, or entirely fabricated responses—a phenomenon commonly referred to as “hallucination” Lee et al. (2018).
@@
To this end, we propose Extract-Refine-Retrieve-Read (ERRR), a simple but effective framework designed for retrieval augmentation systems. The ERRR framework is crafted to bridge the pre-retrieval information gap through tailored query optimization and aims to resolve the inherent limitations of RRR by enabling retrieval based on the specific information needs of the LLM reader. Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer which refines user queries. This refinement either complements or validates the extracted parametric knowledge, ensuring that only essential information is retrieved for generating accurate responses, and minimizing the retrieval of extraneous information that could degrade the quality of the output.
@@ 2.1 Retrieval-Augmented Generation
The integration of retrieval modules to access relevant contextual knowledge has played a crucial role in enhancing Large Language Models (LLMs) in recent years. Initially designed for early sequence-to-sequence models, the Retrieval-Augmented Generation (RAG) framework proposed by Piktus et al. Lewis et al. (2020) has gained substantial traction in the era of LLM. This approach has diversified into a broad array of methods, with ongoing efforts aimed at further augmenting its capabilities. Earlier exploration primarily focused on improving key components, such as upgrading to more powerful pre-trained language models like BERT Devlin et al. (2019) as readers or employing advanced dense retrievers for retrieval tasks Karpukhin et al. (2020). These retrievers encode documents and inputs into dense vectors, facilitating retrieval based on the similarity between the input and retrieved passages.
Recent studies have shifted focus beyond merely enhancing the retriever or reader components, emphasizing the refinement of pre-retrieval and post-retrieval processes. To address the pre-retrieval gap—the disparity between the information retrievable from original queries and the knowledge required for optimal responses—GenRead Yu et al. (2023) replaces the retrieval module with a knowledgeable LLM, thereby narrowing the gap between the user query and retrieval process. It prompts the LLM to generate contextual information for the query, using these generated documents as retrieval results to formulate the final answer. Self-ask Press et al. (2023) proposes an iterative approach using chain-of-thought prompting to generate self-posed questions that refine the response. For the post-retrieval gap—the challenge of creating optimal responses from given information—strategies include document re-ranking or summarization. For instance, PRCA Yang et al. (2023) trains a contextual adaptor module to summarize retrieved documents with a black-box LLM reader.
Several studies have also proposed significant modifications to the original RAG pipeline, introducing complex systems that include both pre-retrieval and post-retrieval modules Rackauckas (2024), and adapting the pipeline into iterative or recursive frameworks Yao et al. (2022); Asai et al. (2023). While these advanced systems demonstrate notable performance enhancements, they incur substantial costs and typically require multiple interactions with LLM. In contrast, our work focuses on refining the single-turn RAG framework, introducing a flexible and trainable module adaptable to existing systems.
@@ 2.2 Query Optimization for Retrieval Augmentation
Recent research highlights a significant discrepancy between input queries and LLM readers for RAG systems, especially under the current trend of using off-the-shelf web search tools or black-box LLMs that are difficult to customize Ma et al. (2023). Typically, these input queries often originate directly from users or specific datasets, which could be either poorly formulated or adhere to a static query format. To overcome these challenges, an effective approach is to optimize the query in the pre-retrieval phase, thereby improving the quality of retrieved information and response generation. The Rewrite-Retrieve-Read (RRR) framework, for instance, trains a query rewriting module using an LLM to better align retrieval queries with LLM readers Ma et al. (2023) that generate the final response, as illustrated in Figure 1. Additionally, RRR introduces a trainable scheme that employs reinforcement learning with Proximal Policy Optimization to fine-tune a small open-source model based on feedback from the LLM reader, achieving improved results. HyDE addresses the demand for accurate information retrieval by creating hypothetical documents and encoding them through unsupervised contrastive learning for efficient retrieval operations Gao et al. (2023). Furthermore, Step-Back Prompting Zheng et al. (2024) converts original queries into high-level abstract questions, aiding LLMs in generating better responses for complex queries requiring abstract thinking.
@@ 3 Methodology
Figure 1: Overview of Extract-Refine-Retrieve-Read (ERRR). Extract-Refine-Retrieve-Read leverages parametric knowledge of LLMs and utilizes a specialized query optimizer to retrieve the knowledge that better aligns with LLM’s needs.
In this section, we elaborate on the details of Extract-Refine-Retrieve-Read (ERRR), a framework for improving the retrieval-augmented LLMs through query optimization for parametric knowledge refinement. Section 3.1 formally defines the central task addressed by ERRR and introduces its key concepts. The design of the framework is discussed in Section 3.2, where we outline a frozen scheme using a black-box LLM reader and standard web search tools. Additionally, Section 3.3 discusses a trainable scheme of the framework.
@@ 3.2 Extract-Refine-Retrieve-Read
Retrieval  To illustrate the adaptability of our module across various retrieval systems and data sources, we utilize two types of retrievers: a black-boxed web search tool and a local dense retrieval system, which are then combined with the original query for processing by the LLM reader.
Generation  We employ an LLM reader to generate the final answer using both the retrieved documents and the original query. Our prompting strategy involves straightforward instruction followed by 1-3 few-shot examples for question answering. These examples are consistently used within each dataset but vary across different datasets to maintain control over the task-specific output format from the LLM reader—for instance, the responses are expected to be concise in certain QA tasks, usually only one or a few words.
@@ 3.3 Trainable Scheme
Given that many powerful LLMs operate as black-box systems, significant challenges such as high computational costs, customization limitations, copyright issues, and connectivity problems have arisen. To address these issues, alongside the conventional frozen scheme, we propose a trainable scheme for our pipeline. Specifically, we fine-tuned a smaller, trainable model utilizing knowledge distillation from a high-performing teacher LLM, leveraging its broadly trained outputs as a good starting point and learning template, and intensively training student models on a distillation dataset of QA questions and generated responses to learn the intricate nuances of query optimization. This streamlined model is then integrated into our pipeline to fulfill the role of query rewriting, originally handled by a frozen LLM.
@@ 4.3 Implementation Details
For all baselines, we utilized GPT-3.5-Turbo as the primary LLM and adhered to their implementation from the original paper. GPT-3.5 Turbo was chosen for its balance of performance and cost, aligning with our focus on optimizing retrieval-augmented generation systems rather than benchmarking generative models themselves. While GPT-4 offers improved capabilities, our emphasis remained on augmenting the model’s utility through query optimization. Notably, for the Trainable RRR, we employed the supervised fine-tuned T5 model checkpoint as the base model. This checkpoint, open-sourced by the original authors, has been warmed up and fine-tuned on multiple datasets to function as the query rewriter. Then we replicated their reinforcement learning process since we replaced the original search tool with the Brave Search Engine. These trainings were conducted on the first 1000 data points for each dataset evaluated, with The training parameters set as follows: a learning rate of 2e-5, 3 epochs, and a batch size of 8.
@@ 4.4 Result
The impact of the ERRR framework is more pronounced in web search retrieval systems, as evidenced by the greater performance enhancement observed in Table 2 compared to dense retrievers in Table 3. This is likely due to the higher quality and broader knowledge span of web-based retrieval systems compared to the static 2018 Wikipedia corpus used for dense retrieval. Notably, the results show that both Frozen RRR and Trainable RRR underperform the Direct method in the PopQA and HotPotQA datasets when using dense retrieval. This underperformance can be attributed to the low-quality results retrieved from the outdated and limited corpus, which includes only Wikipedia passages of constrained length and scope. These limitations lead to an increased retrieval of irrelevant documents, distracting the Large Language Model (LLM) from answering questions correctly.
@@ 4.5 Cost and Latency
Given our method’s emphasis on a conventional single-turn pipeline, it demonstrates superior performance in terms of cost and latency when compared to certain advanced and iterative RAG frameworks. To underscore the cost-efficiency and flexibility of our approach, we conducted a comparative analysis with ReAct Yao et al. (2022) and Self-RAG Asai et al. (2023). This experiment was carried out on 200 randomly selected questions from HotPotQA. The results presented in Table 4 highlight that while still maintaining commendable performance, Frozen ERRR exhibits significantly lower costs, faster processing times, and greater efficiency than other iterative frameworks. Moreover, Trainable ERRR has the potential to further reduce costs, particularly for large datasets, by leveraging an already fine-tuned query optimizer, thereby saving on an additional LLM call to GPT-3.5-Turbo.
@@ 5 Conclusion
In this paper, we present Extract-Refine-Retrieve-Read (ERRR) framework for Retrieval-Augmented Generation (RAG) systems. The ERRR framework is designed to optimize queries, aligning them closely with the specific informational needs of large language models (LLMs) to enhance retrieval augmentation effectiveness. Our experimental results demonstrate that our method surpasses both the naive LLM and native query rewriting framework Rewrite-Retrieve-Read on benchmark datasets such as AmbigQA Min et al. (2020), PopQA Mallen et al. (2022) and HotPotQA Yang et al. (2018), utilizing both web search tools and a dense retriever with local static corpus. It demonstrated ERRR’s remarkable adaptability across a variety of settings, data sources, and retrieval systems. This flexibility ensures that ERRR can be effectively implemented in diverse operational environments, making it a potential and adaptable component for inclusion in more advanced RAG systems. Additionally, we have developed and implemented a trainable scheme for the ERRR framework. This approach is both cost-effective and efficient as it relies on only a fine-tuned T5 model trained on a moderately sized dataset and surpasses the performance of the frozen GPT-3.5-Turbo.
@@ 6 Limitation
We acknowledge that we recognize the existence of more sophisticated Retrieval-Augmented Generation (RAG) approaches such as Self-RAGAsai et al. (2023) and CRAGYan et al. (2024). These advanced systems typically require iterative invocations of the entire pipeline to refine their answers, resulting in exceptionally high computational demands. Due to computational constraints within our study, we focused solely on scenarios that operate in a single-turn manner, wherein each module is invoked only once.
Additionally, our model does not employ any reinforcement learning techniques to enhance the performance of the supervised fine-tuned model. This decision was driven by resource limitations and observed sub-optimal performance when training with a small portion of the dataset using Proximal Policy Optimization (PPO) Schulman et al. (2017), which constrained the potential upper limit of our model’s performance.
For the future development of this work, while the ERRR framework addresses the pre-retrieval gap problem, future work could extend to methods that bridge the post-retrieval gap or incorporate ERRR into more advanced and modular RAG systems to further enhance performance in question-answering tasks. Furthermore, exploring new Reinforcement Learning (RL) algorithms to improve the query optimizer’s performance for specialized tasks is also a possible direction for further exploration.
@@ Appendix A Case Study
To explicitly and intuitively demonstrate the effectiveness of the ERRR compared to the RRR framework, we present two examples in Table 5 comparing their rewritten queries and final outputs. In the first example, the original question is Stories USA starred which actor and comedian from "The Office"?. The query rewriter in RRR framework produces a simplified query, actor comedian "The Office" Stories USA cast, which merely reformulates the original question for clearer web searching. In contrast, the ERRR not only answers correctly in the Parametric Knowledge Extraction phase but also generates refined queries such as "actor and comedian from "The Office" in Stories USA" and "Steve Carell role in Stories USA", These queries not only attempt to validate the actor name of the T̈he Officeb̈ut also attempt to validate the name Steve Carell from the parametric knowledge, enabling the retriever to source better results.
In the second example, the rewritten query from RRR, Pakistani actor writer Islamabad Coke Kahani 2012, rewrites into only a few random keywords from the original question, which fails to facilitate a high-quality search. On the other hand, the first rewritten query from ERRR, Pakistani actor and writer from Islamabad who helped write for Coke Kahani, provides a clearer and more comprehensible query for search possibly inspired by the contextual information from the extracted parametric knowledge. The second rewritten query, Faisal Rehman contributions to Coke Kahani aims to verify the name derived from parametric knowledge, specifically Faisal Rehman. Interestingly, even though the name is incorrect, the information retrieved subsequently clarifies that Faisal Rehman is not the correct actor and writer, which effectively rectifies the LLM’s output. Together with the information gathered from the first query, this leads to a correct final answer. This example illustrates that even if the pseudo-contextual document contains inaccuracies, the ERRR framework, by concentrating on the specific needs of the LLM reader, can still retrieve the most useful information for the LLM reader which results in a correct outcome.

### v4

@@ Abstract
Youan Cong Pritom Saha Akash Cheng Wang Kevin Chen-Chuan Chang University of Illinois Urbana-Champaign, USA {youanc2, pakash2, chengw4, kcchang}@illinois.edu
@@ 1 Introduction
The field of natural language processing (NLP) has witnessed transformative advancements in recent years, largely driven by the advent of Large Language Models (LLMs). These models, trained on vast corpora, have demonstrated exceptional capabilities in understanding human text and generating high-quality responses Kaplan et al. (2020); Clark et al. (2022). They have also proven practical and scalable for various downstream NLP tasks, such as conversational response generation, text summarization, and content recommendation, even in few-shot or zero-shot settings Wu et al. (2023). Despite their strengths, a key limitation of LLMs lies in their reliance on static training data, which causes them to struggle with dynamic or less commonly known information outside their initial training scope. This limitation often leads to outdated, inaccurate, or entirely fabricated responses—a phenomenon commonly referred to as “hallucination” Lee et al. (2018).
@@
To this end, we propose Extract-Refine-Retrieve-Read (ERRR), a straightforward yet effective framework designed for retrieval augmentation systems. The ERRR framework is crafted to bridge the pre-retrieval information gap through tailored query optimization and aims to resolve the inherent limitations of RRR by enabling retrieval based on the specific information needs of the LLM reader. Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer that refines user queries. This refinement either complements or validates the extracted parametric knowledge, ensuring that only essential information is retrieved for generating accurate responses, and minimizing the retrieval of extraneous information that could degrade output quality.
@@ 2.1 Retrieval-Augmented Generation
The integration of retrieval modules to access relevant contextual knowledge has played a crucial role in enhancing Large Language Models (LLMs) in recent years. Initially designed for early sequence-to-sequence models, the Retrieval-Augmented Generation (RAG) framework proposed by Lewis et al. Lewis et al. (2020) has gained substantial traction in the era of LLMs. This approach has diversified into a broad array of methods, with ongoing efforts aimed at further augmenting its capabilities. Earlier exploration primarily focused on improving key components, such as upgrading to more powerful pre-trained language models like BERT Devlin et al. (2019) as readers or employing advanced dense retrievers for retrieval tasks Karpukhin et al. (2020). These retrievers encode documents and inputs into dense vectors, facilitating retrieval based on the similarity between the input and retrieved passages.
Recent studies have shifted focus beyond merely enhancing the retriever or reader components, emphasizing the refinement of pre-retrieval and post-retrieval processes. To address the pre-retrieval gap—the disparity between the information retrievable from original queries and the knowledge required for optimal responses—GenRead Yu et al. (2023) replaces the retrieval module with a knowledgeable LLM, thereby narrowing the gap between the user query and retrieval process. It prompts the LLM to generate contextual information for the query, using these generated documents as retrieval results to formulate the final answer. Self-ask Press et al. (2023) proposes an iterative approach using chain-of-thought prompting to generate self-posed questions that refine the response. For the post-retrieval gap—the challenge of creating optimal responses from given information—strategies include document re-ranking or summarization. For instance, PRCA Yang et al. (2023) trains a contextual adapter module to summarize retrieved documents with a black-box LLM reader.
Several studies have also proposed significant modifications to the original RAG pipeline, introducing complex systems that include both pre-retrieval and post-retrieval modules Rackauckas (2024), and adapting the pipeline into iterative or recursive frameworks Yao et al. (2022); Asai et al. (2023). While these advanced systems demonstrate notable performance enhancements, they incur substantial costs and typically require multiple interactions with LLMs. In contrast, our work focuses on refining the single-turn RAG framework, introducing a flexible and trainable module adaptable to existing systems.
@@ 2.2 Query Optimization for Retrieval Augmentation
Recent research highlights a significant discrepancy between input queries and LLM readers for RAG systems, especially under the current trend of using off-the-shelf web search tools or black-box LLMs that are difficult to customize Ma et al. (2023). Typically, these input queries originate directly from users or specific datasets, which could be either poorly formulated or adhere to a static query format. To overcome these challenges, an effective approach is to optimize the query in the pre-retrieval phase, thereby improving the quality of retrieved information and response generation. The Rewrite-Retrieve-Read (RRR) framework, for instance, trains a query rewriting module using an LLM to better align retrieval queries with LLM readers Ma et al. (2023) that generate the final response, as illustrated in Figure 1. Additionally, RRR introduces a trainable scheme that employs reinforcement learning with Proximal Policy Optimization to fine-tune a small open-source model based on feedback from the LLM reader, achieving improved results. HyDE addresses the demand for accurate information retrieval by creating hypothetical documents and encoding them through unsupervised contrastive learning for efficient retrieval operations Gao et al. (2023). Furthermore, Step-Back Prompting Zheng et al. (2024) converts original queries into high-level abstract questions, aiding LLMs in generating better responses for complex queries requiring abstract thinking.
@@ 3 Methodology
Figure 1: Overview of Extract-Refine-Retrieve-Read (ERRR). ERRR leverages parametric knowledge of LLMs and utilizes a specialized query optimizer to retrieve the knowledge that better aligns with the LLM’s needs.
In this section, we elaborate on the details of Extract-Refine-Retrieve-Read (ERRR), a framework for improving retrieval-augmented LLMs through query optimization for parametric knowledge refinement. Section 3.1 formally defines the central task addressed by ERRR and introduces its key concepts. The design of the framework is discussed in Section 3.2, where we outline a frozen scheme using a black-box LLM reader and standard web search tools. Additionally, Section 3.3 discusses a trainable scheme of the framework.
@@ 3.2 Extract-Refine-Retrieve-Read
Retrieval  To illustrate the adaptability of our module across various retrieval systems and data sources, we utilize two types of retrievers: a black-box web search tool and a local dense retrieval system, which are then combined with the original query for processing by the LLM reader.
Generation  We employ an LLM reader to generate the final answer using both the retrieved documents and the original query. Our prompting strategy involves straightforward instructions followed by 1-3 few-shot examples for question answering. These examples are consistently used within each dataset but vary across different datasets to maintain control over the task-specific output format from the LLM reader—for instance, the responses are expected to be concise in certain QA tasks, usually only one or a few words.
@@ 3.3 Trainable Scheme
Given that many powerful LLMs operate as black-box systems, significant challenges such as high computational costs, customization limitations, copyright issues, and connectivity problems have arisen. To address these issues, alongside the conventional frozen scheme, we propose a trainable scheme for our pipeline. Specifically, we fine-tuned a smaller, trainable model using knowledge distillation from a high-performing teacher LLM. We leveraged the teacher’s outputs as a strong learning template, intensively training the student model on a distillation dataset of QA questions and generated responses to learn the intricate nuances of query optimization. This streamlined model is then integrated into our pipeline to fulfill the role of query rewriting, originally handled by a frozen LLM.
@@ 4.3 Implementation Details
For all baselines, we utilized GPT-3.5-Turbo as the primary LLM and followed the implementation details from their respective original papers. GPT-3.5 Turbo was chosen for its balance of performance and cost, aligning with our focus on optimizing retrieval-augmented generation systems rather than benchmarking generative models themselves. While GPT-4 offers improved capabilities, our emphasis remained on augmenting the model’s utility through query optimization. Notably, for the Trainable RRR, we employed the supervised fine-tuned T5 model checkpoint as the base model. This checkpoint, open-sourced by the original authors, has been warmed up and fine-tuned on multiple datasets to function as the query rewriter. Then we replicated their reinforcement learning process since we replaced the original search tool with the Brave Search Engine. This training was conducted on the first 1000 data points for each dataset evaluated, with The training parameters set as follows: a learning rate of 2e-5, 3 epochs, and a batch size of 8.
@@ 4.4 Result
The impact of the ERRR framework is more pronounced in web search retrieval systems, as evidenced by the greater performance enhancement observed in Table 2 compared to dense retrievers in Table 3. This is likely due to the higher quality and broader knowledge span of web-based retrieval systems compared to the static 2018 Wikipedia corpus used for dense retrieval. Notably, the results show that both Frozen RRR and Trainable RRR underperform the Direct method in the PopQA and HotPotQA datasets when using dense retrieval. This underperformance can be attributed to the low-quality results retrieved from the outdated and limited corpus, which includes only Wikipedia passages of constrained length and scope. These limitations lead to an increased retrieval of irrelevant documents, hindering the LLM from answering questions correctly.
@@ 4.5 Cost and Latency
Given our method’s emphasis on a conventional single-turn pipeline, it demonstrates superior performance in terms of cost and latency when compared to certain advanced and iterative RAG frameworks. To underscore the cost-efficiency and flexibility of our approach, we conducted a comparative analysis with ReAct Yao et al. (2022) and Self-RAG Asai et al. (2023). This experiment was carried out on 200 randomly selected questions from HotPotQA. The results presented in Table 4 highlight that while still maintaining commendable performance, Frozen ERRR exhibits significantly lower costs, faster processing times, and greater efficiency than other iterative frameworks. Moreover, Trainable ERRR has the potential to further reduce costs, particularly for large datasets, by leveraging an already fine-tuned query optimizer, thereby eliminating an additional LLM call to GPT-3.5-Turbo.
@@ 5 Conclusion
In this paper, we present Extract-Refine-Retrieve-Read (ERRR) framework for Retrieval-Augmented Generation (RAG) systems. The ERRR framework is designed to optimize queries, aligning them closely with the specific informational needs of large language models (LLMs) to enhance retrieval augmentation effectiveness. Our experimental results demonstrate that our method surpasses both the naive LLM and native query rewriting framework Rewrite-Retrieve-Read on benchmark datasets such as AmbigQA Min et al. (2020), PopQA Mallen et al. (2022) and HotPotQA Yang et al. (2018), utilizing both web search tools and a dense retriever with local static corpus. These results demonstrated ERRR’s remarkable adaptability across a variety of settings, data sources, and retrieval systems. This flexibility ensures that ERRR can be effectively implemented in diverse operational environments, making it a potential and adaptable component for inclusion in more advanced RAG systems. Additionally, we have developed and implemented a trainable scheme for the ERRR framework. This approach is both cost-effective and efficient as it relies on only a fine-tuned T5 model trained on a moderately sized dataset and surpasses the performance of the frozen GPT-3.5-Turbo.
@@ 6 Limitation
We acknowledge the existence of more sophisticated Retrieval-Augmented Generation (RAG) systems such as Self-RAG Asai et al. (2023) and CRAG Yan et al. (2024). These advanced systems typically require iterative invocations of the entire pipeline to refine their answers, resulting in exceptionally high computational demands. Due to these computational constraints, our study focused solely on scenarios that operate in a single-turn manner, wherein each module is invoked only once. Additionally, our trainable query optimizer does not employ any reinforcement learning (RL) techniques to further refine its alignment with the reader, a decision driven by resource constraints and observed performance degradation when training on a small subset of the dataset using Proximal Policy Optimization (PPO) Schulman et al. (2017).
The effectiveness of the ERRR framework is also inherently tied to the capabilities of the underlying LLMs used for knowledge extraction, query optimization, and final generation. The quality of the extracted knowledge and the refined queries are dependent on the base model’s own internal knowledge and reasoning abilities. Consequently, the framework’s overall performance may vary significantly when implemented with different or future LLMs.
These limitations point toward several directions for future work. Beyond addressing the pre-retrieval gap, future research could also tackle post-retrieval challenges by exploring methods to better re-rank, filter, or synthesize retrieved documents before they are passed to the reader. Furthermore, ERRR could be incorporated as a modular component within more advanced, iterative RAG systems to create a more holistic and powerful RAG pipeline. Finally, exploring novel RL algorithms tailored for this framework could lead to further improvements for the trainable optimizer, allowing it to better realize its potential for adaptation.
7 Acknowledgement
This material is based upon work supported by the National Science Foundation IIS 16-19302 and IIS 16-33755, Zhejiang University ZJU Research 083650, IBM-Illinois Center for Cognitive Computing Systems Research (C3SR) and IBM-Illinois Discovery Accelerator Institute (IIDAI), grants from eBay and Microsoft Azure, UIUC OVCR CCIL Planning Grant 434S34, UIUC CSBS Small Grant 434C8U, and UIUC New Frontiers Initiative. Any opinions, findings, conclusions, or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the funding agencies.
@@ Appendix A Case Study
To explicitly and intuitively demonstrate the effectiveness of the ERRR compared to the RRR framework, we present two examples in Table LABEL:case_study comparing their rewritten queries and final outputs. In the first example, the original question is Stories USA starred which actor and comedian from "The Office"?. The query rewriter in RRR framework produces a simplified query, actor comedian "The Office" Stories USA cast, which merely reformulates the original question for clearer web searching. In contrast, the ERRR not only answers correctly in the Parametric Knowledge Extraction phase but also generates refined queries such as "actor and comedian from "The Office" in Stories USA" and "Steve Carell role in Stories USA", These queries not only attempt to validate the actor name of the T̈he Officeb̈ut also attempt to validate the name Steve Carell from the parametric knowledge, enabling the retriever to source better results.
Example 1
Question Stories USA starred which actor and comedian from "The Office"?
Answer Steven John Carell
RRR
Rewritten Query(s) actor comedian "The Office" Stories USA cast
Output Ricky Gervais (incorrect)
ERRR
Parametric Knowledge The model’s internal knowledge correctly identifies Steve Carell as the star of "Stories USA" and an actor from "The Office," providing additional (though unverified) details about his role and career.
Rewritten Query(s) actor and comedian from "The Office" in Stories USA, Steve Carell role in Stories USA
Output Steven John Carell (correct)
Example 2
Question What Pakistani actor and writer from Islamabad helped write for the 2012 Pakistani comedy drama sitcom, "Coke Kahani"?
Answer Yasir Hussain
RRR
Rewritten Query(s) Pakistani actor writer Islamabad Coke Kahani 2012
Output Ali Abbas (incorrect)
ERRR
Parametric Knowledge The model incorrectly identifies Faisal Rehman as a writer for "Coke Kahani," but correctly notes he is a Pakistani actor and writer. This flawed parametric knowledge still provides a useful starting point for query refinement.
Rewritten Query(s) Pakistani actor and writer from Islamabad who helped write for Coke Kahani, Faisal Rehman contributions to Coke Kahani
Output Yasir Hussain (correct)
Table 5: A case study comparing the RRR and ERRR frameworks. For each example, we present the query and output from the baseline RRR, followed by the extracted parametric knowledge, refined queries, and final output from ERRR.
In the second example, the rewritten query from RRR, Pakistani actor writer Islamabad Coke Kahani 2012, rewrites into only a few random keywords from the original question, which fails to facilitate a high-quality search. On the other hand, the first rewritten query from ERRR, Pakistani actor and writer from Islamabad who helped write for Coke Kahani, provides a clearer and more comprehensible query for search possibly inspired by the contextual information from the extracted parametric knowledge. The second rewritten query, Faisal Rehman contributions to Coke Kahani aims to verify the name derived from parametric knowledge, specifically Faisal Rehman. Interestingly, even though the name is incorrect, the information retrieved subsequently clarifies that Faisal Rehman is not the correct actor and writer, which effectively rectifies the LLM’s output. Together with the information gathered from the first query, this leads to a correct final answer. This example illustrates that even if the pseudo-contextual document contains inaccuracies, the ERRR framework, by concentrating on the specific needs of the LLM reader, can still retrieve the useful information, resulting in a correct outcome.